OpenStudy (anonymous):

Math Analysis: (tan^2x/secx+1) +1

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

i got tan^2x=1-sec^2

OpenStudy (anonymous):

and secx=1/cosx

OpenStudy (anonymous):

I seriously don't get this.. :|

OpenStudy (kinggeorge):

Write sec(x) as 1/cos(x) Write tan(x) as sin(x)/cos(x) See if that helps you.

OpenStudy (kinggeorge):

Also on the bottom, write sec(x)+1 as 1/cos(x) + cos(x)/cos(x) = (1 + cos(x))/cos(x))

OpenStudy (kinggeorge):

\[\sec(x)+1={1 \over \cos(x)} + {\cos(x) \over \cos(x)}\]

OpenStudy (kinggeorge):

\[={{1+\cos(x)}\over \cos(x)}\]From there you it might be more clear on how to simplify farther.